# MASQUE: Localized Adversarial Makeup Generation with Text-Guided Diffusion Models

### v2

@@ Abstract
As face recognition is increasingly adopted for government and commercial services, its potential misuse has raised serious concerns about privacy and civil rights. To counteract, various anti-facial recognition techniques have been proposed for privacy protection by adversarially perturbing face images, among which generative makeup-based approaches are the most popular. However, these methods, designed primarily to impersonate specific target identities, can only achieve weak dodging success rates while increasing the risk of targeted abuse. In addition, they often introduce global visual artifacts or a lack of adaptability to accommodate diverse makeup prompts, compromising user satisfaction. To address the above limitations, we develop MASQUE, a novel diffusion-based framework that generates localized adversarial makeups guided by user-defined text prompts. Built upon precise null-text inversion, customized cross-attention fusion with masking, and a pairwise adversarial guidance mechanism using images of the same individual, MASQUE achieves robust dodging performance without requiring any external identity. Comprehensive evaluations on open-source facial recognition models and commercial APIs demonstrate that MASQUE significantly improves dodging success rates over all baselines, along with higher perceptual fidelity preservation, stronger adaptability to various makeup prompts, and robustness to image transformations.
@@ 1 Introduction
Facial recognition (FR) systems have been adopted in a wide range of security, biometrics, and commercial applications (Parkhi et al., 2015). However, their unregulated deployment poses privacy risks, allowing for unauthorized surveillance and malicious tracking. To address these concerns, anti-facial recognition (AFR) technologies have emerged to protect user privacy from unauthorized FR systems (Wenger et al., 2023). AFR techniques vary depending on which stage of facial recognition they disrupt and typically work by modifying images before they are shared online. Among them, adversarial methods are particularly effective, subtly altering face images to evade detection while preserving their natural appearance. Traditional adversarial approaches, such as noise-based methods (Joon Oh et al., 2017; Yang et al., 2021; Zhong & Deng, 2022), obscure facial features with norm-bounded global perturbations, while patch-based techniques (Xiao et al., 2021; Komkov & Petiushko, 2021) optimize adversarial patterns in localized image regions. Unfortunately, these methods often introduce noticeable visual artifacts, compromising the usability of the produced face images.
To overcome these limitations, recent work has shifted toward generative approaches, leveraging generative adversarial networks (GANs) or diffusion models for anti-facial recognition. Instead of relying on pixel-wise perturbations, these methods generate unrestricted yet semantically meaningful modifications that blend naturally into facial features. A notable direction is makeup-based AFR (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024; Fan et al., 2025), which seamlessly integrates adversarial perturbations into makeup—a plausible approach, as makeup is inherently associated with facial appearance (see Section 2 for detailed discussions of related work).
While makeup-based AFR techniques offer a promising balance between privacy protection and aesthetics, they often struggle to preserve fine-grained facial details or fully adhere to user instructions from diverse prompts. In addition, these methods require images of an external target identity to guide the generation process for adversarial makeup transfer and primarily focus on the impersonation setting for evaluation, increasing the risks of targeted abuse. When considering the more privacy-critical dodging scenarios, their performance in protection success rates often drops significantly.
@@
Inoffensive Identity Protection. Built upon pairwise adversarial guidance, MASQUE ensures high dodging success rates without requiring face images of any external identity other than the victim’s. Such a property reduces the risks of targeted misuse and ethical concerns (Section 3.3).
@@
Through comprehensive experiments on two popular image benchmarks, a range of FR models and their variants, along with multiple makeup text prompts, MASQUE not only shows significantly higher dodging success rates over baseline methods, but is also capable of preserving visual quality and fine details of original face images (Section 5). In addition, we also demonstrate the robustness of MASQUE’s performance to diverse makeup prompts and various image transformations (Section 6). All of the above suggest the potential of MASQUE as a superior solution to real-world AFR applications.
@@ 2 Related Work
Anti-Facial Recognition. Earlier works adopted obfuscation techniques to obscure the facial identity features (Newton et al., 2005) or craft perturbations to fool FR models (Yang et al., 2021). While effective, these methods often compromise image quality, limiting their real-world applicability. Poisoning-based approaches (Cherepanova et al., 2021; Shan et al., 2020) introduce a new protection scheme by injecting subtle adversarial noise into images to degrade the effectiveness of recognition models. These techniques excel at disrupting model training without visibly altering the image, but rely on strong assumptions about when and how the unauthorized FR model is constructed. Recently, adversarial makeup (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024; Fan et al., 2025) has emerged as a promising solution to realize the goal of facial privacy protection. These methods leverage the strong generative capability of generative models to embed adversarial perturbations into natural makeup-based facial modifications, deceiving attackers’ facial recognition models while preserving the aesthetic appeal. For example, Hu et al. (2022) proposed AMT-GAN, which introduces a regularization module and a joint training pipeline for adversarial makeup transfer within the GAN framework (Goodfellow et al., 2014). Advancements in generative models have enhanced the performance of adversarial makeup transfer, such as Clip2Protect (C2P) (Shamshad et al., 2023) and DiffAM (Sun et al., 2024), which adopt a StyleGAN model or a diffusion-based framework, to enable seamless adversarial face modifications with much improved visual quality. In this work, we build upon these developments by leveraging diffusion models to generate visually consistent, localized adversarial makeup for facial privacy protection under dodging scenarios.
@@
More recently, a line of research (Liu et al., 2023; 2024; Wang et al., 2025; Han et al., 2025; Salar et al., 2025) proposed leveraging diffusion models to generate imperceptible AFR-protections without focusing on makeups. For instance, DiffProtect (Liu et al., 2023) uses the diffusion autoencoder model and adversarially guides the semantic code of the original image towards a target face, whereas Salar et al. (2025) improved the prior approach by learning unconditional embeddings as null-text guidance and adversarially modifying the latent code in the latent diffusion model. Our work complements these methods, where we focus on improving the state-of-the-art makeup-based AFR approaches without relying on images from any external identity, while preserving the visual quality of original face images with a strong emphasis on localizing and controlling the added perturbations.
@@
Diffusion-Based Image Editing. Localized editing methods extend diffusion models for targeted modifications. For example, mask-based approaches (Avrahami et al., 2023; Couairon et al., 2022) constrain edits to specific regions using spatial masks, preserving unedited areas but often struggling with structural consistency. Attention-based approaches (Hertz et al., 2022; Tumanyan et al., 2023) guide edits via attention injection, achieving better global structure preservation but suffering from editing leakage. In addition, text-guided diffusion models, capable of synthesizing high-quality images from natural language descriptions, significantly promote the popularization of generative AI. Built on denoising diffusion probabilistic models (Ho et al., 2020), they iteratively refine random noise into coherent images guided by text prompts. Recent advancements in diffusion models have expanded their capabilities to tasks such as localized editing and controllable generation. Mask-based approaches (Avrahami et al., 2023; Couairon et al., 2022) achieve local text-guided modifications by incorporating user-defined constraints, like masks, to confine edits to specific regions. While effective in preserving unedited areas, these methods often struggle to maintain structural consistency within the edited regions, especially in complex scenarios. On the other hand, mask-free attention-based methods (Hertz et al., 2022; Cao et al., 2023; Tumanyan et al., 2023) use attention injection mechanisms to guide edits without requiring explicit masks. These methods excel at preserving the global structure of the image but are prone to editing leakage, where changes unintentionally affect areas beyond the intended region. In this work, we leverage mask-based cross-attention guidance (Mao et al., 2023) to achieve localized adversarial perturbations in the form of makeup. This approach ensures the precise application of modifications to desired regions, guided by user-specified prompts, while addressing the limitations of structural inconsistencies seen in previous methods.
3 Problem Formulation
3.1 Facial Recognition (FR)
@@
We consider an adversarial problem setup, where an attacker adopts unauthorized facial recognition models to identify benign users from their publicly shared facial images. This enables the extraction of sensitive information, posing significant privacy risks. Since many well-trained FR models are readily available—either as open-source implementations or through commercial APIs—attackers can easily obtain automated FR tools to achieve this malicious objective. Formally, let be the input space of face images and be the set of possible identities. Given a collection of online-scraped face images (corresponding to multiple identities), the adversary aims to correctly identify victim users from their facial images as many as possible: , where stands for a face image, is the corresponding ground-truth identity, and denotes a FR model.
Typically, a FR model consists of three components: a feature extractor, a gallery database, and a query matching scheme (see Wenger et al. (2023) for detailed descriptions). Since FR models can vary in design, the attacker may utilize different feature extractors and query-matching schemes. To comprehensively assess performance under black-box conditions, we evaluate a range of well-trained FR models, from open-source models to commercial APIs. In addition, we consider face verification as a specialized FR variant, where the gallery database contains only a single user’s images.
3.2 Anti-Facial Recognition (AFR)
@@
where denotes the AFR perturbation function, is a similarity metric that measures the visual distortion of the AFR-perturbed image with reference to the original image , and is a threshold parameter reflecting the distortion upper bound that can be tolerated.
@@
3.3 Why Dodging instead of Impersonation?
In contrast to the untargeted dodging objective, several state-of-the-art studies (Yang et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024) have focused on improving AFR performance in targeted impersonation scenarios. Impersonation-oriented methods aim to adversarially modify a user’s facial images to mislead a FR model into recognizing them as a specific target identity. Let be the perturbation function with respect to an impersonation-oriented method. In prior work, the impersonation success rate (ISR) is usually computed using the following metric:
where is the target identity with . Below, we explain why impersonation is not appropriate.
From the defender’s perspective, impersonation and dodging serve different goals, as reflected in their distinct evaluation metrics (ISR and DSR). As shown by Zhou et al. (2024), high ISR does not guarantee high DSR. If the target identity is not included in the attacker’s model (e.g., using facial verification with only a single image of the victim identity ), optimizing for higher ISR offers little benefit due to misaligned objectives. As will be shown in Section 1, prior impersonation-oriented AFR methods can only achieve weak protection performance when evaluated in dodging scenarios. In addition, impersonation techniques can lead to a higher likelihood of targeted misuse. Unlike dodging, impersonation-oriented approaches allow users to control the target identity they are mistaken for, making them more susceptible to abuse. This ability can be exploited for malicious purposes, such as falsely associating one’s actions with another individual. For instance, if a user intentionally applies an AFR tool to impersonate a target and engages in illegal activity, FR systems may misidentify as the suspect. This deliberate identity manipulation poses significant ethical and legal risks (Wenger et al., 2023). Therefore, effective privacy protection should aim for untargeted dodging success.
Instead of redirecting recognition to a specifically chosen identity, dodging causes misclassification in an uncontrolled manner, typically as an unknown or incorrect identity. This lack of control makes dodging significantly less susceptible to intentional misuse, as users cannot deliberately impersonate or frame a different party. While dodging still prevents correct identification, it does so without assigning a specific false identity, reducing the risk of targeted deception or malicious intent.
4 Proposed Method: MASQUE
4.1 Limitations of Existing AFR Methods
@@
We conduct preliminary experiments to understand the performance of existing AFR methods in terms of privacy protection and visual quality. Figure 1 shows qualitative comparisons of three makeup styles (“red lipstick”, “blue eyebrow”, and “pink eyeshadow”), evaluating several adversarial-based AFR methods, including one of the best current noise-based methods, TIP-IM (Yang et al., 2021), and three state-of-the-art makeup-based approaches (Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024). For each method, we report both the metrics of impersonation and doding success rates across four black-box facial verification models averaged across randomly selected images from CelebA-HQ (Karras et al., 2017) and the three makeup styles. To ensure a fair comparison with image-guided AFR methods, we either select the corresponding makeup image from a benchmark facial makeup dataset (Li et al., 2018) or generate one by inpainting the makeup on a non-makeup image (see Section 5.1 and Appendix B for more experimental details of this preliminary study).
Weak Protection under Dodging. As discussed in Section 3.3, an ideal AFR should account for dodging scenarios to ensure privacy without dependence on a target identity. However, despite its ethical advantages, dodging remains an underexplored and underdeveloped area in makeup-based AFR. Existing methods that perform well under impersonation settings often exhibit a significant drop in DSR, as shown in Figure 1. This highlights a fundamental gap: current techniques are not optimized for dodging, leaving privacy protection less reliable when a user seeks to avoid recognition rather than impersonate another identity. In addition, while some AFR methods like Shamshad et al. (2023); Fan et al. (2025) introduce a variant for dodging, they still rely on a target identity. This reliance contradicts the very goal of dodging-based AFR, which should function without any identity substitution. Thus, a key challenge in AFR research is developing a robust dodging-based approach that provides strong protection while avoiding the ethical risks associated with impersonation.
Non-Localized Face Modification. Noise-based approaches (TIP-IM) generate global perturbations, which fail to localize the edits. Makeup-based AFR embeds perturbations within makeup to achieve a more natural disguise, but many GAN-based approaches (AdvMakeup, AMT-GAN, and C2P) introduce global artifacts that alter non-facial regions. This lack of precise localization is particularly problematic when users seek to protect only their facial identity while preserving the background. As shown in Figure 1, this issue is especially pronounced in complex backgrounds, where perturbations should remain confined to facial regions but instead extend beyond. In contrast, DiffAM enables better localization, effectively applying makeup perturbations from a reference image. However, its reliance on reference images limits flexibility, particularly when suitable references are unavailable. This highlights the need for methods that achieve both precise localization and greater flexibility.
Limited Prompt-Following Ability. Makeup-based AFR aim methods aim to modify facial appearance to prevent accurate identification by FR models. However, existing methods offer limited user control, making customization and usability a challenge. Below, we elaborate on this challenge for makeup-based AFR methods based on reference images and text prompts, respectively:
Image-reference-based AFR transfers makeup styles from a reference image. While they aim for accurate replication, they struggle with reliability and flexibility. As shown in Figure 1, early GAN-based methods like AMT-GAN fail to transfer makeup consistently, while DiffAM, despite improvements, remains restricted. It only applies makeup to three fixed regions—skin, mouth, and eyes—preventing transfer to other areas like colored eyebrows. In addition, these methods require separate fine-tuning for each reference image, incurring high computational overhead.
Text-prompt-based AFR allows users to describe makeup styles in natural language, eliminating the need for reference images or separate fine-tuning while enhancing privacy. However, this method lacks fine-grained control, as C2P relies solely on CLIP directional loss, often leading to unintended modifications beyond the intended areas.
4.2 Detailed Design

### v3

# Controllable Adversarial Makeup for Privacy via Text-Guided Diffusion

@@ Abstract
As face recognition becomes more widespread in government and commercial services, its potential misuse raises serious concerns about privacy and civil rights. To counteract this threat, various anti-facial recognition techniques have been proposed, which protect privacy by adversarially perturbing face images. Among these, generative makeup-based approaches are the most widely studied. However, these methods, designed primarily to impersonate specific target identities, can only achieve weak dodging success rates while increasing the risk of targeted abuse. In addition, they often introduce global visual artifacts or a lack of adaptability to accommodate diverse makeup prompts, compromising user satisfaction. To address the above limitations, we develop MASQUE, a novel diffusion-based framework that generates localized adversarial makeups guided by user-defined text prompts. Built upon precise null-text inversion, customized cross-attention fusion with masking, and a pairwise adversarial guidance mechanism using images of the same individual, MASQUE achieves robust dodging performance without requiring any external identity. Comprehensive evaluations on open-source facial recognition models and commercial APIs demonstrate that MASQUE significantly improves dodging success rates over all baselines, along with higher perceptual fidelity preservation, stronger adaptability to various makeup prompts, and robustness to image transformations.
@@ 1 Introduction
Facial recognition (FR) systems have been adopted in a wide range of security, biometrics, and commercial applications (Parkhi et al., 2015). However, their unregulated deployment poses privacy risks, allowing unauthorized surveillance and malicious tracking. To address these concerns, anti-facial recognition (AFR) technologies have emerged to protect user privacy from unauthorized FR systems (Wenger et al., 2023). AFR techniques vary depending on which stage of facial recognition they disrupt and typically work by modifying images before they are shared online. Among them, adversarial methods are particularly effective, subtly altering face images to evade detection while preserving their natural appearance. Traditional adversarial approaches, such as noise-based methods (Joon Oh et al., 2017; Yang et al., 2021; Zhong & Deng, 2022), obscure facial features with norm-bounded global perturbations, while patch-based techniques (Xiao et al., 2021; Komkov & Petiushko, 2021) optimize adversarial patterns in localized image regions. Unfortunately, these methods often introduce noticeable visual artifacts, compromising the usability of the produced face images.
To overcome these limitations, recent work has shifted toward generative approaches, leveraging generative adversarial networks (GANs) or diffusion models for anti-facial recognition. Instead of relying on pixel-wise perturbations, these methods generate unrestricted yet semantically meaningful modifications that blend naturally into facial features. A notable direction is the makeup-based AFR (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024; Fan et al., 2025), which seamlessly integrates adversarial perturbations into makeup—a plausible approach, as makeup is inherently associated with facial appearance (see Section 2 for detailed discussions of related work).
While makeup-based AFR techniques offer a promising balance between privacy protection and aesthetics, they often struggle to preserve fine-grained facial details or fully adhere to user instructions from diverse prompts. In addition, these methods require images of an external target identity to guide the generation process for adversarial makeup transfer and primarily focus on the impersonation setting for evaluation, increasing the risks of targeted abuse. When considering more privacy-critical dodging scenarios, their performance in protection success rates often drops significantly.
| Dodging | External ID | Image Quality | Localization | Guidance | Prompt Following | |
| TIP-IM (Yang et al., 2021) | ✗ | ✓ | Low | ✗ | - | - |
| AMT-GAN (Hu et al., 2022) | ✗ | ✓ | Low | ✗ | Image | Low |
| C2P (Shamshad et al., 2023) | ✓ | ✓ | Medium | ✗ | Text | Medium |
| DiffAM (Sun et al., 2024) | ✗ | ✓ | High | ✓ | Image | Medium |
| FPP (Salar et al., 2025) | ✗ | ✓ | High | ✗ | - | - |
| MASQUE (ours) | ✓ | ✗ | High | ✓ | Text | High |
@@
Inoffensive Identity Protection. Built upon pairwise adversarial guidance, MASQUE ensures high dodging success rates without requiring face images of any external identity other than the victim’s. Such a property reduces the risks of targeted misuse and ethical concerns.
@@
Through comprehensive experiments across FR models and makeup text prompts, MASQUE not only achieves significantly higher dodging success rates over baseline methods, but is also capable of preserving the visual quality and fine details of the original face images (Section 5). We also show that MASQUE is robust to more diverse makeup prompts and various image transformations (Section 6). All of the above suggest the potential of MASQUE as a superior solution to real-world AFR applications.
@@ 2 Related Work
Anti-Facial Recognition. Earlier works adopted obfuscation techniques to obscure the facial identity features (Newton et al., 2005) or craft perturbations to fool FR models (Yang et al., 2021). While effective, these methods often compromise image quality, limiting their real-world applicability. Poisoning-based approaches (Cherepanova et al., 2021; Shan et al., 2020) introduce a new protection scheme by injecting subtle adversarial noise into images to degrade the effectiveness of recognition models. These techniques excel at disrupting model training without visibly altering the image but rely on strong assumptions about when and how the unauthorized FR model is constructed. Recently, adversarial makeup (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024; Fan et al., 2025) has emerged as a promising solution to realize the goal of facial privacy protection. These methods leverage the strong generative capability of generative models to embed adversarial perturbations into natural makeup-based facial modifications, deceiving attackers’ facial recognition models while preserving the aesthetic appeal. For example, Hu et al. (2022) proposed AMT-GAN, which introduces a regularization module and a joint training pipeline for adversarial makeup transfer within the GAN framework (Goodfellow et al., 2014). Advancements in generative models have enhanced the performance of adversarial makeup transfer, such as Clip2Protect (C2P) (Shamshad et al., 2023) and DiffAM (Sun et al., 2024), which adopt a StyleGAN model or a diffusion-based framework, to enable seamless adversarial face modifications with much improved visual quality. In this work, we build upon these developments by leveraging diffusion models to generate visually consistent, localized adversarial makeup for facial privacy protection under dodging scenarios.
@@
More recently, a line of research (Liu et al., 2023; 2024; Wang et al., 2025; Han et al., 2025; Salar et al., 2025) proposed leveraging diffusion models to generate imperceptible AFR-protections without focusing on makeup. For instance, DiffProtect (Liu et al., 2023) uses the diffusion autoencoder model and adversarially guides the semantic code of the original image towards a target face, whereas Salar et al. (2025) improved the prior approach by learning unconditional embeddings as null-text guidance and adversarially modifying the latent code in the latent diffusion model. Our work complements these methods, where we focus on improving the state-of-the-art makeup-based AFR approaches without relying on images from any external identity, while preserving the visual quality of original faces with a strong emphasis on localizing and controlling the added perturbations through makeup.
@@
Diffusion-Based Image Editing using Text Guidance. Text-guided diffusion models, capable of synthesizing high-quality images from natural language descriptions, significantly promote the popularization of generative AI. Built on denoising diffusion probabilistic models (Ho et al., 2020), they iteratively refine random noise into coherent images guided by text prompts. Recent advances in diffusion models have expanded their capabilities to tasks such as localized editing and controllable generation. Mask-based approaches (Couairon et al., 2022; Avrahami et al., 2023) achieve local text-guided modifications by incorporating user-defined constraints, like masks, to confine edits to specific regions. While effective in preserving unedited areas, these methods often struggle to maintain structural consistency within the edited regions, especially in complex scenarios. On the other hand, mask-free attention-based methods (Hertz et al., 2022; Cao et al., 2023; Tumanyan et al., 2023) use attention injection mechanisms to guide edits without requiring explicit masks. These methods excel at preserving the global structure of the image but are prone to editing leakage, where changes unintentionally affect areas beyond the intended region. In this work, we leverage mask-based cross-attention guidance (Mao et al., 2023) to achieve localized adversarial perturbations in the form of makeup, ensuring precise modifications to desired regions, guided by user-specified prompts, while addressing the limitations of structural inconsistencies seen in previous methods.
3 Background and Problem Formulation
3.1 Facial Recognition
@@
We consider an adversarial problem setup, where an attacker adopts unauthorized facial recognition models to identify benign users from their publicly shared facial images. This enables the extraction of sensitive information, posing significant privacy risks. Since many well-trained FR models are readily available—either as open-source implementations or through commercial APIs—attackers can easily obtain automated FR tools to achieve this malicious objective. Formally, let be the input space of face images and be the set of possible identities. Given a collection of online-scraped face images (corresponding to multiple identities), the adversary aims to correctly identify victim users from their facial images as many as possible: , where stands for a face image, is the corresponding ground-truth identity, and denotes a FR model. A typical FR model consists of a feature extractor, a gallery database, and a query matching scheme (see Wenger et al. (2023) for detailed descriptions). Since FR models can vary in design, the attacker may utilize different feature extractors and query-matching schemes. To comprehensively assess performance under black-box conditions, we evaluate a range of well-trained FR models, from open-source models to commercial APIs. In addition, we consider face verification as a specialized FR variant, where the gallery database contains only a single user’s images.
3.2 Anti-Facial Recognition
@@
where denotes the AFR perturbation function, stands for a metric that measures the visual distortion of the AFR-perturbed image with reference to the original image , and is a threshold parameter reflecting the distortion upper bound that can be tolerated.
@@
4 Our Methodology
4.1 Limitations of Prior Work
@@
First, we conduct preliminary experiments to better understand prior AFR methods in terms of privacy protection and visual quality. Figure 1 shows qualitative comparisons of three makeup styles (“red lipstick”, “blue eyebrow”, and “pink eyeshadow”), evaluating several adversarial-based AFR methods, including one of the best current noise-based methods, TIP-IM (Yang et al., 2021), three state-of-the-art makeup-based approaches (Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024), and most recent non-makeup diffusion-based approach (Salar et al., 2025). For each method, we report both the metrics of impersonation and dodging success rates across four black-box facial verification models averaged across randomly selected images from CelebA-HQ (Karras et al., 2017) and the three makeup styles. To ensure a fair comparison with image-guided makeup-based AFR methods, we either select the corresponding makeup image from a benchmark facial makeup dataset (Li et al., 2018) or generate one by inpainting the makeup on a non-makeup image (see Section 5.1 and Appendix B for more experimental details of this preliminary study).
Weak Protection under Dodging. In contrast to the untargeted dodging objective, most prior work on anti-facial recognition (Yang et al., 2021; Hu et al., 2022; Shamshad et al., 2023; Sun et al., 2024; Salar et al., 2025) has focused on targeted impersonation, where a user’s facial image is adversarially modified to be recognized as a specific target identity (). These methods are typically evaluated by the impersonation success rate (ISR), which is defined as follows:
However, ISR and the DSR reflect different goals. As Zhou et al. (2024) shows, high ISR does not necessarily imply strong dodging performance. When the target identity is absent from the attacker’s model (e.g., simple verification against victim identity ), optimizing for ISR offers little privacy benefit. Moreover, impersonation methods introduce ethical and legal risks, as they enable deliberate framing of others (Wenger et al., 2023). By contrast, dodging misleads recognition in an uncontrolled manner—typically to an unknown or incorrect identity—thus preventing identification without enabling targeted misuse. For this reason, effective AFR for privacy should prioritize dodging.
More concretely, existing AFR techniques that excel in impersonation settings often suffer a sharp performance drop in DSR (Figure 1), exposing a fundamental gap: current methods are not optimized for dodging, leaving privacy protection unreliable when the user seeks only to avoid recognition. Even works that claim to support dodging (Shamshad et al., 2023; Fan et al., 2025) often depend on a target identity, contradicting the very goal of identity-free protection. Developing robust dodging-based AFR remains a key challenge for enabling reliable and ethically sound privacy defenses.
Limits in Spatial Precision & User Control. Noise-based approaches like TIP-IM produce global pixel-wise perturbations that are often visually perceptible, while GAN-based methods, such as AMT-GAN and C2P, introduce artifacts extending beyond the face area. This lack of spatial precision is problematic when users wish to disguise only their facial identity while preserving natural backgrounds, as shown in Figure 1. DiffAM improves localization by applying makeup perturbations from reference images, but its reliance on such references restricts flexibility when suitable examples are unavailable. Beyond localization, current methods also suffer from limited prompt-following ability, constraining user control and customization. Reference-based AFR struggles with consistent style transfer: early GAN methods (e.g., AMT-GAN) fail to replicate makeup reliably, while DiffAM remains restricted to three fixed facial regions (skin, mouth, eyes), excluding areas like eyebrows, and requires fine-tuning for each reference image. Text-prompt-based AFR removes this dependency and improves usability, but models like C2P, which rely solely on CLIP directional loss, often apply unintended modifications outside the target regions. Together, these issues highlight the need for AFR techniques that are both precisely localized and responsive to user intent.
4.2 Detailed Design of MASQUE